OpenStudy (anonymous):

In the sequence 16, 80, 48, 64, A, B,C,D, each term beyond the second term is a arithmetic mean (average) of the previous two terms. What is the value of D?

OpenStudy (anonymous):

A = (48+64)/2 = 56 B = (64 + 56)/2 = 60 C = (56 + 60)/2 = 58 D = (60 + 58)/2 = 59